i have to find all real solution to equation 5(b-1)^2-4=221 and also i don't know how to apply quadratic formula to this equation.
Answer by stanbon(75887)

i have to find all real solution to equation 5(b-1)^2-4=221 and also i don't know how to apply quadratic formula to this equation.
=======
5(b-1)^2-4=221
5(b-1)^2 = 225
(b-1)^2 = 45
---
(b-1) = +-sqrt(9*5)
---
b-1 = +-3sqrt(5)
---
b = 1+3sqrt(5) or b = 1-3sqrt(5)
================================
Cheers,
Stan H.
========================